A wooden block is placed in a tub of liquid with d
Answer #1

Here ,

let us assume that the density of wooden block is p

Now , as the wooden block is half submerged in the first liquid

relation between density of first liquid and woooden block

p1 = 2*p

p = p1/2

NOw , let the volume of the block is V

for the fraction is x

after pouring the second liquid

Using Archimedes principle

p1 * x * V * g + p2 * (1 - x)*V * g = p * V * g

p1 * x + p2 - p2 * x = p1/2

2 p1 * x + 2 p2 - 2 p2 * x = p1

x = (p1 - 2p2)/(2 *(p1 - p2))

the fraction of the block in the lower liquid is (p1 - 2p2)/(2 *(p1 - p2))